OpenStudy (anonymous):

A rectangle has an area of 14 square centimeters. Its length is five cm greater than its width. Find the dimensions.

OpenStudy (anonymous):

Start off with a good picture, and remember that the area of a rectangle is its length times its width: 

OpenStudy (anonymous):

so 7 and 2?

OpenStudy (anonymous):

Right, w=2. So the width is 2 and the length is w+5=2+5=7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem, glad I could help